A particle of mass 1 gram and charge 5 microcoulomb is thrown at a speed of 20 m per second against a uniform electric field strength 2 x 10 to the power 5 Newton per coulomb the distance travelled by the particle before coming to rest.a)0.1 b)0.2 c)0.3 d)0.4

Grade:12

1 Answers

Khimraj
6 years ago
q = 5*10-6C.
v = 20m/s.
m = 10-3kg.
E= 2*105N/C.
Let distance travelled is d.
Then,
qEd = (½)mv2.
Then d = 0.2m.
Hope it clears. If you like answer then please approve it.